OpenStudy (anonymous):

Write a system of two equations with the variables m and f in each to represent the following problem: To skate at Roller Heaven, each person must pay a monthly membership fee (m) and a fee for each session (f). In January, Art attended 5 sessions and paid $25 in all. In July, he attended 10 sessions, and he paid $40 in all.

OpenStudy (anonymous):

m+fx=#$ Jan: m+5f=25 m+10f=40

OpenStudy (anonymous):

m=25-5f from (1) 25-5f+10f=40 =>f=3 m=10 i think

OpenStudy (anonymous):

Hmm, okay! Thank you.